OpenStudy (anonymous):

Drew saved $325 one year by setting his air conditioner up to 78 degrees from 75 degrees. How much (to the nearest dollar) could he save by setting it up to 77 degrees from 75 degrees?

OpenStudy (anonymous):

@terenzreignz

terenzreignz (terenzreignz):

What's the difference between 78 and 75?

terenzreignz (terenzreignz):

78 - 75 = ?

terenzreignz (terenzreignz):

oh okay... hang on... thinking again :D

OpenStudy (anonymous):

$226.51 $236.61 $236.51

terenzreignz (terenzreignz):

This is trickier than I'm used to...

OpenStudy (anonymous):

@Eyad @msingh @ganeshie8 @Juarismi

OpenStudy (anonymous):

@yrelhan4

OpenStudy (anonymous):

Well, i thin this is way that this is supposed to be solved. So if the increase in cost for 75° is 51% and the increase in the cost of 78° is 18% Then the percentage of saving by setting up to 78 degrees from 75 degrees is 51-18 = 33 And that is $325 Now from 75 from 77 the percentage of saving is 51-28= 23. We're almost done, because, we know that the 33% is $325, and we want to know how many dollars (say x) is the 23% And that hapens to be x= (23*325)/33 And that value matches with one alternative

OpenStudy (primeralph):

Drew is broke lol